OpenStudy (anonymous):

WILL FAN/MEDAL FOR HELP Determine if the lines 9x-2y=16 and 2x-9y=18 are parallel, perpendicular, or neither. a. parallel b. perpendicular c. neither

OpenStudy (anonymous):

The answer is Neither. Neither since these equations don't contain slopes that yield the product of -1 [perpendicular] or are same [parallel]. I hope this helps you!

OpenStudy (anonymous):

C

OpenStudy (anonymous):

It did! Thank you!:) Could you help me with two more?

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

1. Write the equation of the line parallel to x=-16 and passes through the point (9,-10) a. y=-6 b. y=-10 c. x=9 d. x=-10 2. Write the equation of the line perpendicular to 9x+2y=18 and passes through point (1,4). a. 9x-2y=1 b. 9x+2y=17 c. 2x-9=-34 d. 2x+9y=38

OpenStudy (anonymous):

@DavidVines96 ?